40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Entry  Fri Dec 20 21:58:14 2019, Koji, Update, PD QE, PD TEC driver / A2P6 aligned / Lens moved P_20191220_192440_vHDR_On.jpgP_20191220_180929_vHDR_On.jpg
    Reply  Fri Dec 20 22:26:19 2019, Koji, Update, PD QE, PD TEC driver / A2P6 aligned / Lens moved 6x
Message ID: 2482     Entry time: Fri Dec 20 21:58:14 2019     Reply to this: 2483
Author: Koji 
Type: Update 
Category: PD QE 
Subject: PD TEC driver / A2P6 aligned / Lens moved 

== Currently, A2P6 is aligned ==


1) I've brought another TEC driver fro the PD temp control. This unit was borrowed from the 2um ECDL setup. Eventually, we need to return this to ECDL. (Attachment 1)
The PID loop of the TEC control works. But it is not well optimized yet. If you change the target temp too quickly, the TEC out seemed oscillating. Watch the TEC out carefully and change the temp setpoint slowly.
So far I have tried to cool the thermister up to 30kOhm (~232K) and I_TEC was 0.33A.
I did not try further. I felt it was better to cool the PD base for further trial.

2) A part of the alignment study, the beam is aligned to A2P6. Also, the lens position was investigated, and I decided to move the lens ~1 inch away from the window.  (Attachment 2)
    In fact, this allowed us to insert the power meter between the lens and the window. 

Attachment 1: P_20191220_192440_vHDR_On.jpg  1.115 MB  | Hide | Hide all
P_20191220_192440_vHDR_On.jpg
Attachment 2: P_20191220_180929_vHDR_On.jpg  406 kB  | Hide | Hide all
P_20191220_180929_vHDR_On.jpg
ELOG V3.1.3-